# Artificial Intelligence-derived Vascular Age from Photoplethysmography: A Novel Digital Biomarker for Cardiovascular Health

### v3

@@ 2.1 Deep learning model for vascular age prediction
Model 1, adjusted by age, sex, ethnicity and BMI | ||||||
MACCE | 1.02 (1.02-1.03) | <0.005 | 0.79 (0.77-0.82) | <0.005 | 2.37 (2.17-2.59) | <0.005 |
Diabetes | 1.04 (1.04-1.05) | <0.005 | 0.68 (0.64-0.72) | <0.005 | 2.69 (2.42-2.99) | <0.005 |
Hypertension | 1.05 (1.05-1.05) | <0.005 | 0.62 (0.59-0.64) | <0.005 | 2.88 (2.66-3.11) | <0.005 |
Coronary heart disease | 1.03 (1.03-1.04) | <0.005 | 0.71 (0.68-0.74) | <0.005 | 2.20 (2.00-2.42) | <0.005 |
Heart failure | 1.02 (1.02-1.02) | <0.005 | 0.82 (0.77-0.87) | <0.005 | 2.15 (1.82-2.54) | <0.005 |
Myocardial infraction | 1.04 (1.03-1.04) | <0.005 | 0.69 (0.64-0.73) | <0.005 | 2.51 (2.17-2.91) | <0.005 |
Stroke | 1.03 (1.02-1.03) | <0.005 | 0.78 (0.72-0.85) | <0.005 | 2.55 (2.11-3.08) | <0.005 |
All-cause mortality | 1.02 (1.02-1.02) | <0.005 | 0.86 (0.82-0.90) | <0.005 | 2.51 (2.22-2.84) | <0.005 |
@@
Model 2, adjusted by age, sex, ethnicity, BMI, smoking status, hypertension, diabetes, dyslipidemia, and CKD | ||||||
MACCE | 1.02 (1.02-1.02) | <0.005 | 0.82 (0.79-0.85) | <0.005 | 2.24 (2.05-2.45) | <0.005 |
Coronary heart disease | 1.03 (1.03-1.03) | <0.005 | 0.72 (0.69-0.76) | <0.005 | 2.13 (1.94-2.35) | <0.005 |
Heart failure | 1.02 (1.01-1.02) | <0.005 | 0.84 (0.79-0.89) | <0.005 | 2.08 (1.76-2.46) | <0.005 |
Myocardial infraction | 1.03 (1.03-1.04) | <0.005 | 0.70 (0.66-0.75) | <0.005 | 2.41 (2.08-2.79) | <0.005 |
Stroke | 1.02 (1.02-1.03) | <0.005 | 0.80 (0.74-0.87) | <0.005 | 2.41 (2.00-2.91) | <0.005 |
All-cause mortality | 1.02 (1.01-1.02) | <0.005 | 0.89 (0.85-0.94) | <0.005 | 2.35 (2.07-2.66) | <0.005 |
@@
2.3 Vascular age gap classification over time based on serial PPG data is associated with future risk of major adverse cardiovascular and cerebrovascular events
@@
The results based on serial PPG measurements at two time points are shown in Fig. 5. Group G4 (both measurements classified as correct prediction or underestimation) served as the reference. The KM curves and corresponding HRs for future MACCE are presented for Group G1 (both measurements classified as overestimation), Group G2 (first measurement classified as underestimation or correct prediction, second as overestimation), and Group G3 (first measurement classified as overestimation, second as underestimation or correct prediction). The HRs were 5.53 (95% CI: 2.73–11.23, p < 0.005) for Group G1, 3.63 (95% CI: 1.86–7.06, p < 0.005) for Group G2, and 3.02 (95% CI: 2.21–4.13, p < 0.005) for Group G3. The log-rank test p-values for comparisons of G1, G2, and G3 with the reference group G4 were all < 0.001, indicating significant differences. Notably, the HRs were consistently higher than 1 and and showed an increasing trend from Group G3 to Group G1, suggesting that changes in vascular age gap classification over time may provide additional value beyond single-time-point assessments.
@@
2.4 Higher vascular age gap correlates with increased in-hospital mortality
@@
In the external validation using the MIMIC hospital mortality set, we evaluated the association between vascular age gap and in-hospital mortality. Vascular age gap was analyzed both as a continuous variable and a categorical variable (Table 4). For the categorical analysis, underestimation and overestimation were defined using a threshold of 15 years, based on the standard deviation of the overall prediction error.
@@
After adjusting for age, sex, and ethnic background, vascular age gap as a continuous variable was significantly associated with in-hospital mortality, with an odds ratio (OR) of 1.02 per one-year increase (p = 0.01). When treated as a categorical variable, underestimation showed no significant difference compared to the correct prediction group (OR: 0.91, p = 0.49). In contrast, overestimation was associated with a significantly higher risk of in-hospital mortality (OR: 1.86, p < 0.005). Further adjustments for smoking status, hypertension, diabetes, dyslipidemia, and CKD revealed that each additional year of vascular age gap remained marginally associated with increased in-hospital mortality (OR: 1.01, p = 0.06). The underestimation group continued to show no significant difference from the correct prediction group (OR: 0.95, p = 0.71), whereas overestimation remained significantly associated with a higher risk of in-hospital mortality (OR: 1.74, p = 0.01). These findings suggest that a greater vascular age gap, particularly in cases of overestimation, is linked to an elevated risk of in-hospital mortality.
Vascular age gap | Underestimation | Overestimation | ||||
@@
Model 1, adjusted by age, sex, and ethnicity | ||||||
In-hospital mortality | 1.02 (1.00-1.03) | 0.01 | 0.91 (0.69-1.19) | 0.49 | 1.86 (1.24-2.79) | <0.005 |
Model 2, adjusted by age, sex, ethnicity, smoking status, hypertension, diabetes, dyslipidemia, and CKD | ||||||
@@
2.5 AI-vascular age model demonstrates consistent focus on key PPG features across age groups
@@
Figure 6 presents the average PPG waveforms and corresponding saliency maps for three distinct age groups (40, 55, and 70 years old). In the saliency maps, regions with higher brightness indicate areas that attract more attention from the model and are therefore more influential in the prediction process. The saliency maps show that, across all age groups, the model consistently focuses on the region surrounding the diastolic peak, with secondary attention directed towards the systolic peak. This suggests that the model relies on these two key features in the PPG waveform for vascular age prediction. Correspondingly, the average PPG waveforms exhibit age-related changes, particularly in the diastolic and systolic peaks. As age increases, the distance between the two peaks narrows, and the prominence of the diastolic peak gradually diminishes. These observed trends in the PPG waveforms closely match the areas of attention highlighted in the saliency maps, which is consistent with findings in existing literature Yousef et al. (2012); Lin et al. (2023). This alignment further supports the model’s ability to identify features related to vascular aging in the PPG signals, contributing to accurate vascular age estimation.
3 Methods
3.1 Datasets
UK Biobank cohort
The UKB is a large-scale biomedical resource comprising de-identified genetic, lifestyle, and health data, along with biological samples, from over 500,000 participants across the United Kingdom Sudlow et al. (2015). The cohort includes four assessment instances: the initial assessment visit (2006–2010) with 502,150 participants, the first repeat assessment visit (2012–2013) with 20,337 participants, the imaging visit (2014–present) with 90,683 participants, and the first repeat imaging visit (2019–present) with 12,734 participants. Longitudinal follow-up data are also available, enabling comprehensive tracking of participants’ health over time. In this study, PPG signals were extracted from UKB field 4205, and corresponding age information was obtained from UKB field 21003. The PPG signals were recorded using the PulseTrace PCA2 device, with each signal preprocessed and provided by UKB in a standardized format containing 100 sample points per waveform.
PulseDB Dataset
PulseDB is a curated dataset designed specifically for benchmarking cuff-less blood pressure estimation methods Wang et al. (2023). It contains high-quality 10-second PPG segments sampled at 125 Hz from 5,361 subjects, accompanied by demographic information such as age. The dataset consists of two primary subsets: the MIMIC-III subset (n=2,423) and the VitalDB subset (n=2,938). The MIMIC-III subset includes PPG data from patients admitted to the critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012 Johnson et al. (2016a), while the VitalDB subset comprises PPG recordings from surgical patients undergoing routine or emergency procedures at Seoul National University Hospital, Republic of Korea Lee et al. (2022). In this study, the VitalDB subset was used to fine-tune the vascular age prediction model initially developed on the UKB dataset. This fine-tuning step was necessary to account for potential temporal and population-level differences between the UKB and PulseDB datasets, ensuring better generalization and adaptability of the model to new data sources., while the MIMIC-III subset was employed for external validation. Furthermore, the MIMIC-III subset within PulseDB was linked to the MIMIC-III Clinical Database Johnson et al. (2016b) to obtain corresponding in-hospital mortality data for additional analysis.
3.2 Model development
3.2.1 Model architecture
The vascular age prediction model is based on Net1D Hong et al. (2020), a one-dimensional convolutional neural network (CNN) built on a ResNet He et al. (2016) variant. It is selected for its ability to capture subtle and discriminative patterns in PPG signals related to vascular aging. Net1D employs sequential convolutional blocks with residual connections He et al. (2016) and squeeze-and-excitation (SE) modules Hu et al. (2018) to enhance feature extraction. The residual connections mitigate model degradation and stabilize training, while the SE modules adaptively recalibrate channel-wise feature responses to emphasize relevant vascular aging features. This architecture ensures robust and efficient processing of PPG signals for vascular age prediction.
3.2.2 Imbalanced regression in vascular age prediction
In vascular age prediction using the UKB dataset, a significant challenge arises from the highly imbalanced age distribution. Most samples are concentrated within the age range of 50 (second decile) to 69 (ninth decile) years, while the tail regions (37–49 and 70–87 years) contain relatively few samples. Conventional loss functions, such as the MAE, treat all samples equally. Consequently, models trained with these loss functions tend to focus excessively on the majority age group. This causes samples in sparsely populated regions to be biased toward the dense central region, resulting in a pronounced deviation between the distribution of model predictions and the true label distribution. To address this issue, we propose Dist Loss, a distribution-aware loss function that enhances performance in underrepresented regions without sacrificing overall accuracy. The central idea of Dist Loss is to align the distribution of model predictions with the true label distribution, thereby ensuring accurate predictions across all age groups (Figure 1).
The technical details of Dist Loss are described below. Dist Loss comprises two main components: the sample-wise MAE and the distributional MAE. The sample-wise MAE is equivalent to the conventional MAE loss and computes the absolute error between each prediction and its corresponding ground truth, ensuring individual sample accuracy. The distributional MAE, on the other hand, quantifies the discrepancy between the distribution of model predictions and the true label distribution. It is optimized by minimizing the MAE between a pseudo-label sequence (constructed based on the label distribution) and a pseudo-prediction sequence (obtained by sorting the model predictions).
Specifically, we first employ kernel density estimation (KDE) Parzen (1962) to obtain a smooth approximation of the label distribution across the dataset. For each label , a probability is estimated to reflect its overall frequency. For a batch of size , the expected frequency of is given by . Since is generally non-integer, we take its floor value . To ensure that the total frequency matches the batch size , we compute the residual , where denotes the number of unique labels. An auxiliary sequence is then defined as
where indexes the position in the sequence. Essentially, this procedure replicates each label according to its adjusted frequency . For instance, if the label set is with adjusted frequencies , the resulting pseudo-label sequence is . The pseudo-prediction sequence is obtained by sorting the model predictions in the batch. To ensure differentiability, we employ the fast differentiable sorting method described in Blondel et al. (2020).
The overall loss function is then defined as
3.2.3 Experimental setup
The model was implemented using the PyTorch 2.2 framework and trained on an NVIDIA GeForce RTX 4090 GPU (24GB memory). The training process spanned 80 epochs, utilizing the Adam optimizer with an initial learning rate of 3e-3. To enhance training stability, we employed a large batch size of 2,048 and applied L2 weight regularization with a factor of to mitigate overfitting. For the Dist loss calculation, Gaussian KDE was used with a bandwidth parameter of to model the age label distribution, which spans 21 to 111 years. Model performance was tracked using the MAE as the primary evaluation metric throughout the training.
3.2.4 Model interpretability
To enhance the interpretability of the model, this study employs saliency maps to visualize the regions of the PPG signals that the model prioritizes when predicting vascular age. Saliency maps are generated by computing the gradients of the input signal with respect to the model’s output, thereby highlighting the model’s sensitivity to various components of the signal during prediction. Specifically, regions exhibiting larger gradient magnitudes are indicative of areas that exert a stronger influence on the model’s prediction. To refine the visual clarity, Gaussian smoothing and interpolation techniques are applied to the saliency maps, minimizing noise and improving the discernibility of relevant signal features.
3.3 Outcome definition
3.4 Serial PPG analysis
Given that some participants in the UKB dataset attended multi-time point assessments, resulting in serial PPG data across different time points, we further investigated the application of AI-vascular age in such serial PPG data. Following the exclusion criteria outlined in Figure 2, all serial PPG data were excluded from the model development set, leaving 58,232 PPG records from 27,108 participants. Our analysis specifically focused on two distinct time points of serial PPG data: for participants with more than two PPG records, we selected their first two PPG data for analysis. In this analysis, we divided the participants into four groups based on the vascular age gap at the two time points: G1 (overestimation at both time points), G2 (underestimation or correct prediction at the first time point and overestimation at the second), G3 (overestimation at the first time point and underestimation or correct prediction at the second), and G4 (underestimation or correct prediction at both time points). We studied the differences in the risk of future incident MACCE across these groups.
3.5 Statistical analysis
To assess the consistency between AI-vascular age and calendar age, we calculated Pearson correlation coefficients and the MAE. We introduced the vascular age gap, defined as the difference between AI-vascular age and calendar age. This metric was then used to evaluate the clinical utility of AI-vascular age. To explore its predictive value, we performed multivariable-adjusted Cox proportional hazards regression analyses. In the UKB cohort, confounding variables were adjusted in two hierarchical models: Model 1 included age, sex, ethnicity, and BMI; Model 2 further adjusted for hypertension, diabetes, dyslipidemia, smoking status, and CKD. For the MIMIC mortality set, adjustments were limited to age, sex, and ethnicity due to data limitations, with BMI excluded due to missing height information in a significant proportion of the patient population.
In the UKB survival analysis, participants with missing BMI data or those lost to follow-up were excluded from the clinical evaluation cohort, resulting in a study population of 111,561 participants and 141,745 PPG records. For censored cases without observed events, follow-up was administratively censored as of December 31, 2022. The log-rank test was employed to compare KM curves.
4 Discussion
@@
In this study, we developed a CNN to predict vascular age directly from PPG signals. The model was trained and evaluated on the UKB dataset and externally validated using the MIMIC subset from the PulseDB dataset. To assess deviations in vascular aging, we introduced the vascular age gap, defined as the difference between AI-predicted vascular age and calendar age, based on a single PPG measurement. Our findings demonstrate that the vascular age gap serves as an independent risk factor for the primary outcome, MACCE, and secondary outcomes such as mortality, HF, MI, stroke, CHD, hypertension, and diabetes within the UKB cohort. External validation using the MIMIC subset further revealed that the vascular age gap is significantly associated with the risk of in-hospital mortality, suggesting its potential clinical application in both the general population and critically ill patients.
@@
We also conducted an analysis of serial PPG measurements within the UKB cohort. Participants were categorized into four groups based on vascular age gap classifications at two time points: G1 (overestimation in both measurements), G2 (not overestimation in the first measurement but overestimation in the second), G3 (overestimation in the first measurement but not in the second), and G4 (not overestimation in either measurement). Compared to G4, participants in the other groups exhibited significantly higher risks for future MACCE events: G1 (HR 5.53, 95% CI: 2.73–11.23, p < 0.005), G2 (HR 3.63, 95% CI: 1.86–7.06, p < 0.005), and G3 (HR 3.02, 95% CI: 2.21–4.13, p < 0.005). The HR showed a progressively increasing trend from G3 to G1. This pattern may reflect the cumulative effect of vascular aging. G1 participants consistently displayed advanced vascular aging across both measurements, signifying prolonged exposure to poor cardiovascular health and the highest overall risk. In G2, the transition to overestimation during the second measurement suggests an accelerated progression of vascular aging, which reflects a recent deterioration in cardiovascular health. Consequently, their risk is elevated but remains lower than that of G1. For G3, although the second measurement indicates improvement in cardiovascular health, the long-term effects of previous excessive vascular aging may persist, leading to a moderately increased risk compared to G4, but still lower than G2.
@@
To address the common challenge of data imbalance in regression tasks, particularly the age imbalance in the UKB dataset, we propose a novel loss function, Dist Loss. This method aligns the distribution of model predictions with the ground truth label distribution, effectively mitigating the risk of over-predicting samples from underrepresented regions into areas with higher label density. Our approach provides a straightforward and effective solution for handling imbalanced data distributions, offering a practical and reliable method for regression tasks under such conditions.
@@
In tackling the interpretability challenges often associated with deep learning models, this study utilizes saliency map visualizations to identify areas of focus within the input signals. The results reveal that the model predominantly concentrates on the region around the diastolic peak of the PPG waveform when predicting vascular age, with secondary emphasis on the systolic peak. These findings are consistent with observed variations in the average PPG waveforms across different age groups. This approach provides a useful framework for understanding and explaining the model’s prediction behavior, supporting conclusions from existing research.
CVDs are a leading global health threat, and individuals’ ability to understand and manage their cardiovascular health is crucial in reducing the societal burden of these diseases. The aging population has contributed to a steady rise in the incidence of cardiovascular conditions, underscoring the increasing need for early screening and diagnosis. The widespread adoption of wearable devices has made PPG signals, due to their non-invasive, convenient, and continuous monitoring capabilities, an increasingly important tool in cardiovascular health management. Research focusing on cardiovascular diseases using PPG signals has gained traction in recent years. Studies have explored the relationship between PPG signals and vascular aging, with some indicating that increased arterial stiffness due to vascular aging can lead to the disappearance of the dicrotic notch and inflection points in PPG waveforms. Most previous research has relied on traditional feature engineering methods to extract parameters such as the reflection index from PPG signals to study their relationship with cardiovascular diseases. However, these manually extracted parameters are not always intuitive, and the reliance on specialized knowledge limits their broader application in personalized cardiovascular health management. Thus, there is a growing need for more intuitive and easily interpretable cardiovascular health indicators for individuals.
Unlike traditional waveform parameters such as pulse wave velocity, age is a highly intuitive and easily understood metric, widely applicable in clinical settings. By leveraging PPG signals, an indicator reflecting vascular aging offers substantial value in practical applications. Building upon this concept, this study proposes using deep learning techniques to predict vascular age directly from PPG signals, thus quantifying the degree of vascular aging. This approach provides a straightforward, actionable, and easily understandable cardiovascular health indicator. The underlying rationale is that increased arterial stiffness due to vascular aging induces changes in the PPG waveform, embedding information on vascular aging within the signal. Through validation across multiple datasets, we demonstrate the utility of this vascular age indicator in both single and serial PPG measurements. Furthermore, its association with in-hospital mortality highlights its potential value in critically ill patients. This vascular age indicator is easily accessible, as it can be obtained through devices such as smart wristbands. Its intuitive nature further enhances its practical application. In daily life, individuals can use this indicator to gain a better understanding of their cardiovascular health, potentially prompting proactive cardiovascular disease screenings. In clinical settings, this indicator can assist healthcare professionals in evaluating a patient’s hemodynamic status, guiding intervention strategies to improve patient outcomes.
5 Conclusion
This study introduces an accessible and intuitive physiological age indicator that reflects an individual’s cardiovascular health status. Preliminary results support its potential application in both the general population and critically ill patients. Future research will focus on further exploring the clinical value of the AI-based vascular age indicator from several perspectives: 1) Investigating its relationship with specific cardiovascular diseases; 2) Exploring its application in long-term continuous monitoring, leveraging the richer data obtained from such monitoring; and 3) Conducting deeper validation of the indicator’s significance in real-world settings, offering actionable health management solutions and potential interventions.
@@ 5 Conclusion
This study used data from the UKB, which is globally accessible to approved researchers (http://ukbiobank.ac.uk/register-apply/). UKB has obtained ethical approval from the North West-Haydock Research Ethics Committee (REC reference: 21 /NW/0157). This study was approved by the UKB under application 90018. The PulseDB dataset is publicly available at https://github.com/pulselabteam/PulseDB. Access to the MIMIC-III Clinical Database requires completion of the Collaborative Institutional Training Initiative course and approval from PhysioNet (https://physionet.org/content/mimiciii/1.4).
The implementation of Dist Loss is available in the GitHub repository at https://github.com/Ngk03/Dist-Loss. The model configuration and pretrained weights can be found at https://huggingface.co/Ngks03/PPG-VascularAge.

### v4

@@ Results
| Model 1, adjusted for age, sex, ethnicity and BMI | ||||||
| MACCE | 1.02 (1.02–1.03) | 2.43 10-100 | 0.79 (0.77–0.82) | 2.48 10-36 | 2.37 (2.17–2.59) | 8.46 10-80 |
| Diabetes | 1.04 (1.04–1.05) | 1.00 10-128 | 0.68 (0.64–0.72) | 1.58 10-39 | 2.69 (2.42–2.99) | 3.14 10-77 |
| Hypertension | 1.05 (1.05–1.05) | 1.00 10-128 | 0.62 (0.59–0.64) | 1.02 10-114 | 2.88 (2.66–3.11) | 1.00 10-128 |
| Coronary heart disease | 1.03 (1.03–1.04) | 1.00 10-128 | 0.71 (0.68–0.74) | 4.36 10-56 | 2.20 (2.00–2.42) | 4.62 10-58 |
| Heart failure | 1.02 (1.02–1.02) | 8.62 10-26 | 0.82 (0.77–0.87) | 4.32 10-11 | 2.15 (1.82–2.54) | 2.28 10-19 |
| Myocardial infarction | 1.04 (1.03–1.04) | 1.26 10-77 | 0.69 (0.64–0.73) | 9.38 10-32 | 2.51 (2.17–2.91) | 5.16 10-35 |
| Stroke | 1.03 (1.02–1.03) | 1.89 10-25 | 0.78 (0.72–0.85) | 6.52 10-10 | 2.55 (2.11–3.08) | 1.53 10-22 |
| All-cause mortality | 1.02 (1.02–1.02) | 2.17 10-36 | 0.86 (0.82–0.90) | 5.47 10-10 | 2.51 (2.22–2.84) | 4.85 10-48 |
@@
| Model 2, adjusted for age, sex, ethnicity, BMI, smoking status, hypertension, diabetes, dyslipidemia, and CKD | ||||||
| MACCE | 1.02 (1.02–1.02) | 3.78 10-79 | 0.82 (0.79–0.85) | 2.36 10-28 | 2.24 (2.05–2.45) | 1.25 10-69 |
| Coronary heart disease | 1.03 (1.03–1.03) | 1.00 10-128 | 0.72 (0.69–0.76) | 1.70 10-49 | 2.13 (1.94–2.35) | 4.57 10-53 |
| Heart failure | 1.02 (1.01–1.02) | 1.84 10-20 | 0.84 (0.79–0.89) | 7.46 10-09 | 2.08 (1.76–2.46) | 1.74 10-17 |
| Myocardial infarction | 1.03 (1.03–1.04) | 5.12 10-68 | 0.70 (0.66–0.75) | 1.56 10-27 | 2.41 (2.08–2.79) | 1.21 10-31 |
| Stroke | 1.02 (1.02–1.03) | 3.47 10-21 | 0.80 (0.74–0.87) | 3.86 10-08 | 2.41 (2.00–2.91) | 6.65 10-20 |
| All-cause mortality | 1.02 (1.01–1.02) | 5.91 10-25 | 0.89 (0.85–0.94) | 3.28 10-06 | 2.35 (2.07–2.66) | 2.80 10-41 |
| Model 3, adjusted for age, sex, SBP, use of antihypertensive treatment, smoking status, diabetes, total cholesterol, and HDL cholesterol. | ||||||
| MACCE | 1.02 (1.02–1.02) | 1.47 10-56 | 0.84 (0.81–0.88) | 2.04 10-15 | 2.37 (2.14–2.62) | 8.71 10-64 |
| Coronary heart disease | 1.03 (1.03–1.04) | 3.07 10-96 | 0.74 (0.71–0.78) | 4.01 10-31 | 2.33 (2.10–2.59) | 1.37 10-55 |
| Heart failure | 1.02 (1.01–1.02) | 3.89 10-18 | 0.86 (0.80–0.92) | 1.08 10-05 | 2.41 (2.01–2.89) | 3.52 10-21 |
| Myocardial infarction | 1.04 (1.03–1.04) | 3.12 10-52 | 0.71 (0.66–0.77) | 5.64 10-19 | 2.62 (2.23–3.08) | 2.44 10-31 |
| Stroke | 1.02 (1.02–1.03) | 3.95 10-13 | 0.85 (0.77–0.93) | 5.10 10-04 | 2.33 (1.87–2.91) | 4.31 10-14 |
| All-cause mortality | 1.01 (1.01–1.02) | 1.38 10-14 | 0.94 (0.89–0.99) | 2.29 10-02 | 2.32 (2.02–2.66) | 7.38 10-33 |
Results in the table were obtained using Cox proportional hazards regression.
@@
AI-PPG age gap classification over time based on serial PPG data is associated with future risk of major adverse cardiovascular and cerebrovascular events
@@
The results based on serial PPG measurements at two time points are shown in Fig. 5. Group G4 (both measurements classified as correct prediction or underestimation) served as the reference. The KM curves and corresponding HRs for future MACCE are presented for Group G1 (both measurements classified as overestimation), Group G2 (first measurement classified as underestimation or correct prediction, second as overestimation), and Group G3 (first measurement classified as overestimation, second as underestimation or correct prediction). The HRs were 5.53 (95% CI: 2.73–11.23, p = 1.5010-6) for Group G1, 3.63 (95% CI: 1.86–7.06, p = 2.6710-4) for Group G2, and 3.02 (95% CI: 2.21–4.13, p = 1.2310-9) for Group G3. The log-rank test p-values for comparisons of G1, G2, and G3 with the reference group G4 were all < 0.001, indicating significant differences. Notably, the HRs were consistently higher than 1 and and showed an increasing trend from Group G3 to Group G1, suggesting that changes in AI-PPG age gap classification over time may provide additional value beyond single-time-point assessments.
@@
Overestimated AI-PPG age correlates with increased in-hospital mortality in critical ill patients
@@
In the external validation using the MIMIC hospital mortality set, we evaluated the association between vascular age gap and in-hospital mortality. AI-PPG age gap was analyzed both as a continuous variable and a categorical variable (Table 4). For the categorical analysis, underestimation and overestimation were defined using a threshold of 15 years, based on the standard deviation of the overall prediction error.
@@
After adjusting for age, sex, and ethnic background, AI-PPG age gap as a continuous variable was significantly associated with in-hospital mortality, with an odds ratio (OR) of 1.02 per one-year increase (p = 0.01). When treated as a categorical variable, underestimation showed no significant difference compared to the correct prediction group (OR: 0.91, p = 0.49). In contrast, overestimation was associated with a significantly higher risk of in-hospital mortality (OR: 1.86, p = 4.0410-4). Further adjustments for smoking status, hypertension, diabetes, dyslipidemia, and CKD revealed that each additional year of AI-PPG age gap remained marginally associated with increased in-hospital mortality (OR: 1.01, p = 0.06). The underestimation group continued to show no significant difference from the correct prediction group (OR: 0.95, p = 0.71), whereas overestimation remained significantly associated with a higher risk of in-hospital mortality (OR: 1.74, p = 0.01). These findings suggest that a greater AI-PPG age gap, particularly in cases of overestimation, is linked to an elevated risk of in-hospital mortality.
| AI-PPG age gap | Underestimation | Overestimation | ||||
@@
| Model 1, adjusted for age, sex, and ethnicity | ||||||
| In-hospital mortality | 1.02 (1.00-1.03) | 0.01 | 0.91 (0.69-1.19) | 0.49 | 1.86 (1.24-2.79) | 4.04 10-4 |
| Model 2, adjusted for age, sex, ethnicity, smoking status, hypertension, diabetes, dyslipidemia, and CKD | ||||||
@@
Results in the table were obtained using Cox proportional hazards regression.
AI-PPG age model demonstrates consistent focus on key PPG features across age groups
@@
Figure 6 presents the average PPG waveforms and corresponding saliency maps for three distinct age groups (40, 55, and 70 years old). In the saliency maps, regions with higher brightness indicate areas that attract more attention from the model and are therefore more influential in the prediction process. The saliency maps show that, across all age groups, the model consistently focuses on the region surrounding the diastolic peak, with secondary attention directed towards the systolic peak. This suggests that the model relies on these two key features in the PPG waveform for age prediction. Correspondingly, the average PPG waveforms exhibit age-related changes, particularly in the diastolic and systolic peaks. As age increases, the distance between the two peaks narrows, and the prominence of the diastolic peak gradually diminishes. These observed trends in the PPG waveforms closely match the areas of attention highlighted in the saliency maps, which is consistent with findings in existing literature 38, 39. This alignment further supports the model’s ability to identify features related to vascular aging in the PPG signals, contributing to accurate age estimation.
Discussion
@@
In this study, we developed a deep learning-based framework to estimate AI-PPG age for streamlined cardiovascular health assessment. The model was trained and evaluated on the UKB dataset and externally validated using the MIMIC subset of the PulseDB dataset. Our results demonstrate that the AI-PPG age gap, defined as the difference between AI-PPG age and chronological age, serves as an independent risk factor for the primary outcome MACCE and secondary outcomes, including mortality, HF, MI, stroke, CHD, hypertension, and diabetes in the UKB cohort. External validation further confirmed its association with in-hospital mortality in the MIMIC subset, underscoring its potential clinical utility in both general and critically ill populations.
@@
We also conducted an analysis of serial PPG measurements within the UKB cohort. Participants were categorized into four groups based on the AI-PPG age gap classifications at two time points: G1 (overestimation in both measurements), G2 (not overestimation in the first measurement but overestimation in the second), G3 (overestimation in the first measurement but not in the second), and G4 (not overestimation in either measurement). Compared to G4, participants in the other groups exhibited significantly higher risks for future MACCE events: G1 (HR 5.53, 95% CI: 2.73–11.23, p = 1.5010-6), G2 (HR 3.63, 95% CI: 1.86–7.06, p = 2.6710-4), and G3 (HR 3.02, 95% CI: 2.21–4.13, p = 1.2310-9). The HR showed a progressively increasing trend from G3 to G1. This pattern may reflect the cumulative effect of vascular aging. G1 participants consistently displayed advanced vascular aging across both measurements, signifying prolonged exposure to poor cardiovascular health and the highest overall risk. In G2, the transition to overestimation during the second measurement suggests an accelerated progression of vascular aging, which reflects a recent deterioration in cardiovascular health. Consequently, their risk is elevated but remains lower than that of G1. For G3, although the second measurement indicates improvement in cardiovascular health, the long-term effects of previous excessive vascular aging may persist, leading to a moderately increased risk compared to G4, but still lower than G2. These results demonstrate the potential of addictive value of serial PPG data. These findings highlight the additive value of serial PPG measurements in capturing dynamic changes in vascular aging and associated cardiovascular risk.
@@
To address the common challenge of data imbalance in regression tasks 40, particularly the age imbalance in the UKB dataset, we propose a loss function, Dist Loss. This method aligns the distribution of model predictions with the ground truth label distribution, effectively mitigating the risk of over-predicting samples from underrepresented regions into areas with higher label density. Our approach provides a straightforward and effective solution for handling imbalanced data distributions, offering a practical and reliable method for regression tasks under such conditions.
@@
In tackling the interpretability challenges often associated with deep learning models, this study utilizes saliency map visualizations to identify areas of focus within the input signals. The results reveal that the model predominantly concentrates on the region around the diastolic peak of the PPG waveform when predicting AI-PPG age, with secondary emphasis on the systolic peak. These findings are consistent with observed variations in the average PPG waveforms across different age groups. This approach provides a useful framework for understanding and explaining the model’s prediction behavior, supporting conclusions from existing research 38, 39.
CVDs remain the leading global health threat, and empowering individuals to understand and manage their cardiovascular health is essential to reducing the associated societal burden 6, 41. The rapid proliferation of wearable devices, particularly those equipped with embedded PPG sensors, offers a promising avenue for non-invasive, convenient, and continuous hemodynamic monitoring 42, 14. In this context, we introduce AI-PPG age, an AI-derived biomarker representing biological age that serves as an intuitive and interpretable proxy for vascular health, promoting individual awareness and self-monitoring. Importantly, AI-PPG age is derived exclusively from raw PPG signals, enabling seamless integration into wearable platforms such as smartwatches. This approach offers two key clinical benefits: first, it supports large-scale cardiovascular screening and risk stratification through wearable devices; second, it facilitates targeted clinical follow-up and personalized management for individuals identified as high risk. By enabling early detection of vascular health changes and guiding timely interventions, this tiered strategy holds significant potential to improve cardiovascular outcomes.
This study has several limitations. First, the results demonstrated a moderate correlation between AI-PPG age and calender age, which may be attributed to the limited age-related information captured in the relatively short PPG recordings. While this does not compromise the utility of AI-PPG age in cardiovascular risk stratification, improving its numerical alignment with calender age could enhance interpretability in practical applications. One potential approach is to apply post hoc correction strategies, such as the bias adjustment methods proposed in previous studies 43, 44, 45. Second, the waveform of PPG signals is not only affected by vascular aging but also by other physiological factors such as blood pressure and the use of vasoactive medications. In this study, to ensure ease of use and convenience, we developed the AI-PPG age prediction model using only raw PPG signals without incorporating additional covariates. While this design improves applicability, it may limit the model’s ability to comprehensively capture cardiovascular health. Future work may explore combining AI-PPG age with other cardiovascular-related indicators at the application stage to further improve assessment performance. Third, following previous studies 46, 47, 23, we used the standard deviation of the AI-PPG age gap as a threshold for stratifying individuals. However, this data-driven threshold may vary across populations, raising concerns about its generalizability. While our findings in the UKB cohort support a 9-year threshold for the general population, further validation is needed before applying this cutoff in other demographic or clinical settings. Finally, AI-PPG age is intended to serve as an auxiliary indicator of cardiovascular risk rather than a standalone diagnostic tool. Its use in real-world settings should be integrated with other clinical information to avoid over-intervention due to outliers or underestimation of risk due to occasional prediction errors.
Despite these limitations, our study demonstrates that AI-PPG age is a scalable and non-invasive digital biomarker with potential for broad implementation in wearable devices. By supporting continuous cardiovascular health monitoring and enabling early risk stratification at the population level, AI-PPG age offers a promising approach to enhance preventive cardiology and inform personalized clinical decision-making.
@@
This study used data from the UK Biobank. Access to UK Biobank data is not publicly available and requires an application. Data are provided to approved researchers worldwide (https://www.ukbiobank.ac.uk/). The PulseDB dataset is publicly available at https://github.com/pulselabteam/PulseDB. Access to the MIMIC-III Clinical Database requires completion of the Collaborative Institutional Training Initiative course and approval from PhysioNet (https://physionet.org/content/mimiciii/1.4).
Code Availability
The implementation of Dist Loss is available in the GitHub repository at https://github.com/Ngk03/Dist-Loss 48. The model configuration and pretrained weights can be found at https://huggingface.co/Ngks03/PPG-VascularAge 49.
Competing interests
The authors declare no competing interests.